# WLM: Dynamics of an isolated Dwarf Irregular Galaxy Under Ram Pressure in the Local Group

### v1

@@ 6.1 Inner Component-1
Gas in the moment-0 map for this component is more centrally packed, both along the major and minor axes. The central iso-velocity contours in the moment-1 map of Component-1 are oriented along a PA of about 165 degrees as seen in the second row of Fig. 10, similar to the moment-1 map from the non-decomposed cube in the first row. In contrast, the inner velocity contours of Component-2 are nearly perpendicular to the PA of WLM. This suggests that we have successfully separated the component that contributes to the 165 degree contours in the global moment-1 map. This off-axis orientation also points towards this component’s bar-like nature. The PV diagram for this component does not show any flattening at the outskirts and shows solid-body like increase in rotational velocity on both the approaching and receding sides all the way till at least 1.7 kpc. However, a careful examination of the PV diagram indicates a very central component, whose orientation is shifted from the overall PV diagram. This suggests component-1 to be more complex than a single bar. We will study the impact of such a structure on the dynamics of a galaxy like WLM interacting with the IGM in a future study with full hydrodynamical modelling. This is also seen in the rotation curve of this component. These bar-like motions contribute to the steep rotation curve in the inner regions. WLM is well known in the literature for a high star formation rate at the center of the galaxy (Albers et al. 2019). Along with the high inclination of WLM, this renders classical tests such as the Tremaine-Weinberg method for bars on a gas velocity field inconclusive (Hernandez et al. 2005). Gaseous bar-like structures have also been reported in several dwarf irregulars such as NGC 6822, NGC 3741, NGC 2915 and DDO 168 (Meurer et al. 1996; Banerjee et al. 2013; Patra and Jog 2019).
Anagha A.G. and Banerjee (2025) have also recently shown that gaseous bars can form and remain stable through hydrodynamical simulations. Another important feature which points towards a central bar-like structure is the distinct two-component split seen in the PV diagram. Bureau and Freeman (1999) show that a separate kinematic component can be a diagnostic for boxy peanut-shaped structures in galaxies. An ancillary support to the component to be a bar-like also comes from its Sersic index of 0.6, as A. Gadotti (2009) suggested that bars usually have Sersic indices between 0.5-1.
@@ 6.2 Outer Component-2
The outer Component-2 in both the morphological and kinematic decomposition shows extended and disk-like features. It seems to trace a more spread-out component of the gas distribution along the major and minor axes of the galaxy. Looking at the velocity field we see that at inner radii, we have a rapid increase in velocity on both sides of the galaxy, which progressively evolve into V-shaped contours often seen in galactic disks. The central iso-velocity contours in the moment-1 map are nearly perpendicular to the PA= 175 degree major axis, indicating clear distinction from the off-axis bar-like Component-1. The dispersion of this component decreases progressively from the center to the outskirts; such steady variation is often seen in galactic disks supporting the validity of our decomposition (Hammer et al. 2017). The classical double-horn profile, which is the kinematic signature of HI rotating disks is seen in Component-2 is seen in Fig. 13. The two horns correspond to the outer parts of the disk where the rotation curve is flat, producing a large contribution at the maximum projected velocities. The relative dip near the systemic velocity arises because gas at central velocities comes from smaller radii and contributes less to the total flux. In addition, the double-horn shape becomes more distinct and stronger as galaxies are viewed closer to edge-on where the observed line-of-sight velocities closely reflect the true rotational velocities (Giovanelli and Haynes 1988). In the PV diagram for this component in Fig. 9 we can clearly see velocities flattening beyond one kiloparsec on the approaching side, which is also seen in its rotation curve in Fig. 11.
@@ 7 Conclusions
WLM is a Local-group gas-rich dwarf irregular whose gas kinematics and dynamics are perturbed by ram pressure effects exerted by the surrounding IGM. Yang et al. (2022) showed that WLM is continuously losing gas as it travels through the IGM. Our revised MeerKAT-16 moment maps show that 13 % of the total gas mass has been presently lost from the main body just in the field of view. Considering that WLM is a gas-rich and star-forming dwarf galaxy at the current epoch, its dynamics was likely dominated by gas motions in the past. Using our new high-resolution MeerKAT 64 observations, we have been able to show that the gas kinematics of WLM beyond 1.5 kiloparsec is severely impacted by its motion through the IGM, consequently affecting the rotation curve and any subsequent mass estimates from it.
Our morphological and kinematic decomposition shows that WLM appears to have an inner bar-like structure which is identified by its solid-body inner rotation curve, which seems more resilient to perturbations, while the disk-like outer component is characterized by a double-horn profile. The latter component is severely impacted by ram pressure effects, which contribute to the rotation curve asymmetry between the receding and the approaching sides.
Dynamical studies of dwarf galaxies often focus on isolated bodies with minimal tidal interactions to better understand their internal structure in order to estimate their total mass. Here we find that dwarf motions with respect to the local IGM can generate ram-pressure effects sufficiently strong to affect the gas dynamics at the outskirts of their disks. It suggests that more consideration must be given to understand the distribution and density of the ionized medium through which galaxies travel.
In the near future, we will present detailed hydrodynamical simulations in an attempt to model WLM and its interaction with the IGM to further probe the above scenario and further probe how ram pressure affects the galaxy differentially. The local group IGM is sparse enough to be undetected by present-day direct observations. Accurately modeling WLM and its interaction with the medium will allow us to put realistic constraints on the local IGM density as well as on the WLM total mass estimate.
@@ Acknowledgements.
- <span style=”font-variant:small-caps;”>sofia</span> 2 – an automated, parallel H <span style=”font-variant:small-caps;”>i</span> source finding pipeline for the WALLABY survey. Monthly Notices of the Royal Astronomical Society 506 (3), pp. 3962–3976 (en). External Links: ISSN 0035-8711, 1365-2966, Link, Document Cited by: §2.1.
- SOFIA 2 - an automated, parallel H I source finding pipeline for the WALLABY survey. MNRAS 506 (3), pp. 3962–3976. External Links: Document, 2106.15789, ADS entry Cited by: §2.2.

### v3

@@ 6.1 Inner Component-1
Gas in the moment-0 map for this component is more centrally packed, along both the major and minor axes. The central iso-velocity contours in the moment-1 map of Component-1 are oriented along a PA of about 165 degrees as seen in the second row of Fig. 10, similar to the moment-1 map from the non-decomposed cube in the first row. In contrast, the inner velocity contours of Component-2 are nearly perpendicular to the PA of WLM. This suggests that we have successfully separated the component that contributes to the 165 degrees contours in the global moment-1 map. This off-axis orientation also points towards this component’s bar-like nature. The PV diagram for this component does not show any flattening at the outskirts and shows solid-body like increase in rotational velocity on both the approaching and receding sides all the way till at least 1.7 kpc. However, a careful examination of the PV diagram indicates a very central component, whose orientation is shifted from the overall PV diagram. This suggests component-1 to be more complex than a single bar. We will study the impact of such a structure on the dynamics of a galaxy like WLM interacting with the IGM in a future study with full hydrodynamical modelling. This is also seen in the rotation curve of this component. These bar-like motions contribute to the steep rotation curve in the inner regions. WLM is well known in the literature for a high star formation rate at the centre of the galaxy (Albers et al. 2019). Along with the high inclination of WLM, this renders classical tests such as the Tremaine-Weinberg method for bars on a gas velocity field inconclusive (Hernandez et al. 2005). Gaseous bar-like structures have also been reported in several dwarf irregulars such as NGC 6822, NGC 3741, NGC 2915 and DDO 168 (Meurer et al. 1996; Banerjee et al. 2013; Patra and Jog 2019).
Anagha A.G. and Banerjee (2025) have also recently shown that gaseous bars can form and remain stable through hydrodynamical simulations. Another important feature that points towards a central bar-like structure is the distinct two-component split seen in the PV diagram. Bureau and Freeman (1999) show that a separate kinematic component can be a diagnostic for boxy peanut-shaped structures in galaxies. An ancillary support to the component to be a bar-like also comes from its Sérsic index of 0.6, as A. Gadotti (2009) suggested that bars usually have Sérsic indices between 0.5 and 1.
@@ 6.2 Outer Component-2
The outer Component-2 in both the morphological and kinematic decomposition shows extended and disc-like features. It seems to trace a more spread-out component of the gas distribution along the major and minor axes of the galaxy. Looking at the velocity field we see that at inner radii, we have a rapid increase in velocity on both sides of the galaxy, which progressively evolve into V-shaped contours often seen in galactic discs. The central iso-velocity contours in the moment-1 map are nearly perpendicular to the PA= 175 degrees major axis, indicating clear distinction from the off-axis bar-like Component-1. The dispersion of this component decreases progressively from the centre to the outskirts; such steady variation is often seen in galactic discs supporting the validity of our decomposition (Hammer et al. 2017). The classical double-horn profile, which is the kinematic signature of HI rotating discs is seen in Component-2 is seen in Fig. 13. The two horns correspond to the outer parts of the disc where the rotation curve is flat, producing a large contribution at the maximum projected velocities. The relative dip near the systemic velocity arises because gas at central velocities comes from smaller radii and contributes less to the total flux. In addition, the double-horn shape becomes more distinct and stronger as galaxies are viewed closer to edge-on where the observed line-of-sight velocities closely reflect the true rotational velocities (Giovanelli and Haynes 1988). In the PV diagram for this component in Fig. 9 we can clearly see velocities flattening beyond one kiloparsec on the approaching side, which is also seen in its rotation curve in Fig. 11.
@@ 7 Conclusions
WLM is a gas-rich LG dwarf irregular whose gas kinematics and dynamics are perturbed by ram pressure effects exerted by the surrounding IGM. Yang et al. (2022) show that WLM is continuously losing gas as it travels through the IGM. Our revised MeerKAT-16 moment maps show that 13 % of the total gas mass has been lost from the main body just in the field of view. Considering that WLM is a gas-rich and star-forming dwarf galaxy at the current epoch, its dynamics was likely dominated by gas motions in the past. Using our new high-resolution MeerKAT-64 observations, we have been able to show that the gas kinematics of WLM beyond 1.5 kiloparsec is severely impacted by its motion through the IGM, which consequently affects the rotation curve and any subsequent mass estimates based on it.
Our morphological and kinematic decomposition shows that WLM appears to have an inner bar-like structure that is identified by its solid-body inner rotation curve; this structure seems more resilient to perturbations. The disc-like outer component is characterised by a double-horn profile and is severely impacted by ram pressure effects, which contribute to the rotation curve asymmetry between the receding and the approaching sides.
Dynamical studies of dwarf galaxies often focus on isolated bodies with minimal tidal interactions to better understand their internal structure and in turn estimate their total mass. Here we find that dwarf motions with respect to the local IGM can generate sufficiently strong ram pressure effects to affect the gas dynamics at the outskirts of their disks. It suggests that more consideration must be given to understanding the distribution and density of the ionised medium through which galaxies travel.
In the near future, we will present detailed hydrodynamical simulations in an attempt to model WLM and its interaction with the IGM to further probe the above scenario and further probe how ram pressure affects the galaxy differentially. The LG IGM is sparse enough to be undetected by present-day direct observations. Accurately modelling WLM and its interaction with the medium will allow us to place realistic constraints on the local IGM density as well as on the WLM’s total mass.
@@ Acknowledgements.
- <span style=”font-variant:small-caps;”>sofia</span> 2 – an automated, parallel H <span style=”font-variant:small-caps;”>i</span> source finding pipeline for the WALLABY survey. Monthly Notices of the Royal Astronomical Society 506 (3), pp. 3962–3976 (en). External Links: ISSN 0035-8711, 1365-2966, Link, Document Cited by: §2.1, §2.2.